# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 3.3 Task incremental scenarios
Task Incremental Scenario (TIS) comes into the picture when we have a multi-task problem and a single adaptive model is desired. Each task is considered an episode. Thus each episode has disjoint label space. However, there is ambiguity in the literature with task incremental scenarios as the other scenarios (class and domain) are frequently referred to as task incremental scenarios by different research communities. For example, Ravishankar et al. (2019) create the first episode having 2 chest x-ray views (2ch vs. PLAX) and then the next episode with 2 other views (4ch vs. PSAX) as a task incremental scenario, which can also be a class incremental scenario. Further, another example in this line can be found in the work by Baweja et al. (2018). Here, the authors create a task incremental scenario with 2 episodes for brain MRI segmentation application where the $1^{st}$ episode was multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the $2nd$ episode was the segmentation of white matter lesions. If we have multiple tasks that are very different, then it would certainly be the task incremental; however, if the tasks are more close then it is merely a design choice whether to treat them as class incremental or task incremental. Another work where Kaustaban et al. (2022) explore domain incremental scenario as organ shift for tumor classification problem where the first episode is a colon cancer dataset with nine classes (CRC dataset), and the next episode is a breast cancer dataset with 2 classes (PatchCam dataset). Similar is the case with Sadafi et al. (2023) where White Blood Cells (WBC) classification from 3 datasets, each having a different number of classes, is regarded as domain incremental, which can also be regarded as task-incremental. Derakhshani et al. (2022) curate class incremental and task incremental scenarios from the same dataset taken from MedMNIST Yang et al. (2023a) where a dataset is divided into non-overlapping classes leading to 4 episodes. If the model is evaluated only on the learnt class in the specific episode, then they regard it as task incremental; otherwise, if evaluated on cumulative classes from all seen episodes, then class incremental scenario.
Kaustaban et al. (2022) say that any of the scenarios among instance, class, and domain incremental can be treated as task incremental, provided that each incoming data stream (episode) is treated as a distinct task. For a comprehensive overview, we have tabulated all the works along with their claimed task incremental scenario in Table 4.
@@
Reference (year) Application (# Episodes) Description
@@
Karani et al. (2018) cross-scanners, hospitals, or acquisition protocols (4 Ep.) MR brain segmentation on various combination using 5 datasets from different source: Human Connectome Project (Van Essen et al. 2013), Alzheimer’s Disease Neuroimaging Initiative (ADNT ), Autism Brain Imaging Data Exchange (Di Martino et al. 2014) and Information eXtraction from Images (IXI )
@@
Lenga et al. (2020) Cross-sites (2 Ep.) Chest x-ray classification on datasets from 2 source: ChestX-ray14 dataset from NIH Clinical Center and MIMIC-CXR dataset from the Beth Israel Deaconess Medical Center
Özgün et al. (2020) Different age ranges, MRI field strengths & presence of pathologies or motion artifacts (total 4 Ep.) Brain MRI segmentation datasets: 4 Ep. were created from 3 datasets: CANDI (Kennedy et al. 2012) (1 Ep.), ADNI (Jack Jr et al. 2008) (1 Ep.), and MALC (Asman and Landman 2013) (2 Ep. based on age)
Hofmanninger et al. (2020a) Total 3 Ep. (a) different scanner parmeter (b) target-shift Chest CT synthetic classification task: (a) first to second episode have change in scanner parameter (Hofmanninger et al. 2020b) (b) target shift is introduced intentionally in the third episode by imprinting a synthetic target structure in the form of a cat on random locations, rotations and varying scale in 50% of the samples.
Morgado et al. (2021) Appearance and body-parts (2 Ep.) Dermatological imaging modality classification (full-body, anatomic, macroscopic, and dermoscopic classes): For each of the class author considered disjoint images in the two episodes, for example, in the second episode, legs and arms images were considered for full-body class, images containing hands or feet only considered for anatomic, images containing face region for macroscopic class, and pink colored images for dermoscopic class.
@@
Zhang et al. (2021) Cross-sites (4-6 Ep.) (a) Prostate segmentation: MRI datasets across 6 sites: RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) optic cup and disc segmentation: public multi-site fundus image datasets from 4 sources (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
@@
Derakhshani et al. (2022) Across organs or modality (4 Ep.) disease classification: 4 epi using 4 dataset ordered as: BloodMNIST, OrganaMNIST, PathMNIST, TissueMNIST (MedMNIST repository)
@@
González et al. (2022) (a) Cross-domain (5 Ep.) (b) manual image contrast change (5 Ep.) MRI hippocampus segmentation: (a) HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019) (b) MRI hippocampus segmentation: data transformation applied using TorchIO library on Decathlon (Simpson et al. 2019) as intensity rescaling, affine transformations, rotation, translation
@@
Kaustaban et al. (2022) (a) Multi-organ (2 Ep.) (b) scanning protocol (5 Ep.) (a) Tumor classification: CRC (a colon cancer dataset with 9 classes), PatchCam (a breast cancer dataset with 2 classes) (b) 5 domain scenario were simulated by changing H&E composition in CRC dataset
Shu et al. (2022) (a) Imaging protocol (2 Ep.) (b) multi-source (3 epi) (a) Low and high-quality fundus images (retinal images) from EyeQ dataset as domain shift condition (Fu et al. 2019), (b) fundus disease classification (AMD, DR, glaucoma, myopia, and normal classes): ODIR (ODIR ), R&R (RIADD (RIADD ) + REFUGE (REFUGE )), and iSee (Fang et al. 2020)
Li et al. (2022) Multi-scanner vendors (4 Ep.) CMR based Cardiac segmentation of left ventricle, right ventricle, and left ventricle myocardium: 4 scanner vendors (Siemens, Philips, General Electric, Cannon) from M&Ms (Campello et al. 2021) dataset
Sadafi et al. (2023) Cross-sites (3 Epi) WBC classification datasets: Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Bera et al. (2023) (a) Cross-center (4 Ep.) (b) cross-center (2 Ep.) (a) Binary prostate segmentation: Prostate158 (Adams et al. 2022) $\rightarrow$ NCI-ISBI (NCI ) $\rightarrow$ Promise12 (Litjens et al. 2014) $\rightarrow$ Decathlon (Antonelli et al. 2022) (b) binary hippocampus segmentation: Drayd (Denovellis et al. 2021) $\rightarrow$ HarP (Boccardi et al. 2015))
Zhu et al. (2023) (a) Cross-site (6 Ep.) (b) cross-site & cross-modality (2 Ep.) (c) same-site & cross-modality (2 Ep.) (a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites (12–30 scans/site) (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015) (b) multi-class (liver, left and right kidneys, and spleen) abdominal segmentation between 30 CT and 20 MRI T2-SPIR scans (c) muscle segmentation of 13 lower-leg muscles and bones between 30 MRI T1 and 30 mDixon scans
@@ 3.4 Domain incremental scenarios
Domain Incremental Scenario (DIS) is the most popular and frequently observed category of CL scenarios for medical applications. This scenario aligns with the idea of learning in a changing environment where the datasets from different domains (e.g., research sites, hospitals, imaging modality, image acquisition protocol, etc.) are incrementally encountered over time, which thus involve covariance shift-induced discrepancies in data.
@@
Yang et al. (2023b) MedMNIST (Yang et al. 2023a) based disease classification (4 Ep.) CIS+DIS: 3 datasets (PathMNIST, DermaMNIST, and OrganAMNIST) as source domain and then 3 more datasets (RetinaMNIST, BreastMNIST, BloodMNIST) as domain shift but classes are added in class incremental fashion with 1 class at a time
Wang et al. (2023b) CIS (a) segmentation for endoscopy (2-3 Ep.) (b) surgical instrument segmentation (2 Ep.) Incrementally adding segmentation structure from EDD2020 (Ali et al. 2021, 2020) dataset which includes 5 classes as Barrett’s esophagus, cancer, high-grade dysplasia, polyp, and suspicious. Initially, 3-4 classes are in 1st episode, and then 2 or 1 classes are incrementally added in subsequent episodes. (b) EndoVis18 (Allan et al. 2020) and EndoVis17 (Allan et al. 2019) datasets as two episodes having some overlapping classes
Liu et al. (2023a) MRI-based brain tumor segmentation (4 Ep.) CIS+DIS Incrementally adding more structures to segmentation task from dataset acquired from different vendors/centers
@@ 3.5 Simulated or Hybrid Settings
It may not be always the case that the model always observes a particular kind of incremental scenario, for example, domain incremental condition. The CL model developed for handling domain incremental conditions may not perform well for a class incremental condition. Thus, there are attempts to evaluate the same model on different CL scenarios; however, a separate evaluation is followed for each kind of incremental CL scenario. In contrast, having a set of episodes with a mix of CL scenarios is close to real-life conditions. Thus it becomes important to also test the performance in hybrid settings. An attempt in direction is made by Sadafi et al. (2023) where other than pure domain incremental and class incremental, a hybrid incremental scenario is also considered. The authors create a sequence of episodes with both class and domain-incremental cases on multi-site WBC classification datasets. They create a long sequence of episodes with a total of 12 episodes. First, they set up a class incremental scenario within the Matek-19 Matek et al. (2019) dataset by adding 3-4 classes in each subsequent episode up to a total of 4 episodes. After this a new domain, i.e., the INT-20 dataset was considered in the same fashion, contributing 4 more episodes. After this, a third domain i.e., a new dataset Acevedo-20 Acevedo et al. (2020) was introduced with 4 episodes in class incremental fashion.
Another novel direction for having a mix of scenarios is considered by Liu et al. (2023a). In contrast to traditional class incremental settings where class-incremental methods are developed without considering data shift, Liu et al. (2023a) incrementally add more structures for brain tumor segmentation application (class incremental) where the incremental structure can be from distinct sites/scanners/MRI modality (domain shift conditions).
@@
Further, there are some attempts to curate shifts in datasets and various incremental scenarios instead of collecting actual datasets reflecting the situation. For example, Kaustaban et al. (2022) simulated domain shift in H&E stain exhibiting distinct appearance caused by different staining protocols. This dataset is designed to simulate real-world scenarios, where data shifts occur due to differences in scanners, stainers, reagents, and other factors. They considered 9 classes from CRC Kather et al. (2019) which is a H&E stain-based colorectal cancer classification dataset. Each class is divided into 5 disjoint sets for 5 episodes. The first set, regarded as Domain1 is without any alteration and the next 4 sets for each class undergo various changes to create domain shift. A short description of 4 domain shifts is as follows: Domain 2 (increased stain intensity, simulating, for example, concentration increase of the eosin and/or hematoxylin solutions, each with a different extent of change). Domain 3 (decreased eosin stain intensity, simulating, for example, slides prepared from many years ago with fading stain). Domain 4 (change of hue, simulating, for example, change of reagent manufacturer, scanner or stainer) Domain 5 (change of hue and saturation, simulating, for example, change of reagent manufacturer, scanner or stainer). finally, authors used this augmented dataset for setting class-incremental, domain-incremental, data-incremental, as well as task-incremental scenarios.
@@
Figure 6: The proposed CL taxonomy categorizes various CL models according to their underlying design principles. The abbreviations used are as follows: ER for Experience-reply, GR for Generative-reply, LR for Latent-reply, PF for Prior-focused, DF for Data-focused, and MBO for Memory Buffer Optimization. 1. Hofmanninger et al. (2020a), 2. Perkonigg et al. (2021), 3. Perkonigg et al. (2022), 4. Karthik et al. (2022), 5. Bera et al. (2023), 6. Li et al. (2022), 7. Srivastava et al. (2021), 8. Baweja et al. (2018), 9. McClure et al. (2018), 10. van Garderen et al. (2019), 11. Özgün et al. (2020), 12. Zhang et al. (2021), 13. Zhang et al. (2023d), 14. Chen and Tang (2022), 15. Zhu et al. (2023), 16. Liu et al. (2023a), 17. Karani et al. (2018), 18. González et al. (2022), 19. Memmel et al. (2021), 20. Bayasi et al. (2021), 21. Ozdemir et al. (2018), 22. Ozdemir and Goksel (2019), 23. Chee et al. (2023), 24. Li et al. (2020), 25. Ravishankar et al. (2019), 26. Liu et al. (2022), 27. Akundi and Sivaswamy (2022), 28. Shu et al. (2022), 29. Zhang et al. (2023b), 30. Sadafi et al. (2023), 31. Yang et al. (2023b),

### v4

@@ 3.3 Task incremental scenarios
Task Incremental Scenario (TIS) comes into the picture when we have a multi-task problem and a single adaptive model is desired. Each task is considered an episode. Thus each episode has disjoint label space. However, there is ambiguity in the literature with task incremental scenarios as the other scenarios (class and domain) are frequently referred to as task incremental scenarios by different research communities. For example, Ravishankar et al. (2019) create the first episode having 2 chest x-ray views (2ch vs. PLAX) and then the next episode with 2 other views (4ch vs. PSAX) as a task incremental scenario, which can also be a class incremental scenario. Further, another example in this line can be found in the work by Baweja et al. (2018). Here, the authors create a task incremental scenario with 2 episodes for brain MRI segmentation application where the $1^{st}$ episode was multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the $2nd$ episode was the segmentation of white matter lesions. If we have multiple tasks that are very different, then it would certainly be the task incremental; however, if the tasks are more close then it is merely a design choice whether to treat them as class incremental or task incremental. Another work where Kaustaban et al. (2022) explore domain incremental scenario as organ shift for tumor classification problem where the first episode is a colon cancer dataset with nine classes (CRC dataset), and the next episode is a breast cancer dataset with 2 classes (PatchCam dataset). Similar is the case with Sadafi et al. (2023) where White Blood Cells (WBC) classification from 3 datasets, each having a different number of classes, is regarded as domain incremental, which can also be regarded as task-incremental. Derakhshani et al. (2022) curate class incremental and task incremental scenarios from the same dataset taken from MedMNIST Yang et al. (2023a) where a dataset is divided into non-overlapping classes leading to 4 episodes. If the model is evaluated only on the learned class in the specific episode, then they regard it as task incremental; otherwise, if evaluated on cumulative classes from all seen episodes, then class incremental scenario. Kaustaban et al. (2022) say that any of the scenarios among instance, class, and domain incremental can be treated as task incremental, provided that each incoming data stream (episode) is treated as a distinct task. For a comprehensive overview, we have tabulated all the works along with their claimed task incremental scenario in Table 6.
@@
Reference (year) Shift source (# Episodes) Description
@@
Karani et al. (2018) Cross-scanners, hospitals, or acquisition protocols (4 Ep.) MR brain segmentation on various combinations using 5 datasets from different sources: Human Connectome Project (Van Essen et al. 2013), Alzheimer’s Disease Neuroimaging Initiative (ADNT ), Autism Brain Imaging Data Exchange (Di Martino et al. 2014) and Information eXtraction from Images (IXI )
@@
Lenga et al. (2020) Cross-sites (2 Ep.) Chest X-ray classification on datasets from 2 sources: ChestX-ray14 dataset from NIH Clinical Center and MIMIC-CXR dataset from the Beth Israel Deaconess Medical Center
Özgün et al. (2020) Different age ranges, MRI field strengths & presence of pathologies or motion artifacts (4 Ep.) Brain MRI segmentation datasets: 4 Ep. were created from 3 datasets: CANDI (Kennedy et al. 2012) (1 Ep.), ADNI (Jack Jr et al. 2008) (1 Ep.), and MALC (Asman and Landman 2013) (2 Ep. based on age)
Hofmanninger et al. (2020a) Different scanner parameters and target-shift (3 Ep.) Chest CT synthetic classification task: the second episode have change in scanner parameter as compared to first episode (Hofmanninger et al. 2020b), then target shift is introduced intentionally in the third episode by imprinting a synthetic target structure in the form of a cat on random locations, rotations and varying scale in 50% of the samples.
Morgado et al. (2021) Appearance and body-parts (2 Ep.) Dermatological imaging modality classification (full-body, anatomic, macroscopic, and dermoscopic classes): For each of the classes, author considered disjoint images in the two episodes, for example, in the second episode, legs and arms images were considered for full-body class, images containing hands or feet only considered for anatomic, images containing face region for macroscopic class, and pink colored images for dermoscopic class.
@@
Zhang et al. (2021) Cross-sites (6 Ep.), (4 Ep.) (a) Prostate segmentation: MRI datasets across 6 sites: RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) optic cup and disc segmentation: public multi-site fundus image datasets from 4 sources (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
@@
Derakhshani et al. (2022) Across organs or modality (4 Ep.) Disease classification: 4 episodes using 4 datasets ordered as: BloodMNIST, OrganaMNIST, PathMNIST, TissueMNIST (MedMNIST repository)
@@
González et al. (2022) (a) Cross-domain (5 Ep.), (b) manual image contrast change (5 Ep.) MRI hippocampus segmentation: (a) HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019), (b) data transformation applied using TorchIO library on Decathlon (Simpson et al. 2019) as intensity rescaling, affine transformations, rotation, translation
@@
Kaustaban et al. (2022) (a) Multi-organ (2 Ep.), (b) scanning protocol (5 Ep.) (a) Tumor classification: CRC (a colon cancer dataset with 9 classes), PatchCam (a breast cancer dataset with 2 classes), (b) Tumor classification: 5 domain shift scenarios were simulated by changing H&E composition in CRC dataset
Shu et al. (2022) (a) Imaging protocol (2 Ep.), (b) multi-source (3 Ep.) (a) Low and high-quality fundus images (retinal images) from EyeQ dataset as domain shift condition (Fu et al. 2019), (b) fundus disease classification (AMD, DR, glaucoma, myopia, and normal classes) across 3 datasets: ODIR (ODIR ), R&R (RIADD (RIADD ) + REFUGE (REFUGE )), and iSee (Fang et al. 2020)
Li et al. (2022) Multi-scanner vendors (4 Ep.) CMR based Cardiac segmentation of left ventricle, right ventricle, and left ventricle myocardium: 4 episodes as 4 scanner vendors (Siemens, Philips, General Electric, Cannon) from M&Ms (Campello et al. 2021) dataset
Sadafi et al. (2023) Cross-sites (3 Ep.) WBC classification across 3 datasets: Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Bera et al. (2023) (a) Cross-center (4 Ep.), (b) cross-center (2 Ep.) (a) Binary prostate segmentation: Prostate158 (Adams et al. 2022) $\rightarrow$ NCI-ISBI (NCI ) $\rightarrow$ Promise12 (Litjens et al. 2014) $\rightarrow$ Decathlon (Antonelli et al. 2022), (b) binary hippocampus segmentation: Drayd (Denovellis et al. 2021) $\rightarrow$ HarP (Boccardi et al. 2015))
Zhu et al. (2023) (a) Cross-site (6 Ep.), (b) cross-site & cross-modality (2 Ep.), (c) same-site & cross-modality (2 Ep.) (a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites (12–30 scans/site) (Liu et al. 2020a; Bloch et al. 2015; Lemaître et al. 2015), (b) multi-class (liver, left and right kidneys, and spleen) abdominal segmentation between 30 CT and 20 MRI T2-SPIR scans, (c) muscle segmentation of 13 lower-leg muscles and bones between 30 MRI T1 and 30 mDixon scans
@@
Table 8: List of various domain shift scenarios in literature (Continued)
Reference (year) Shift source (# Episodes) Description
Byun et al. (2023) Various demographics, collection periods, camera types, and image quality (2-3 Ep.) (a) Diabetic retinopathy severity classification across 2 datasets (Messidor-2, APTOS), (b) dermoscopy skin lesion detection across 3 datasets (BCN2000, PAD-UEFS-20, HAM10000)
Sun et al. (2023b) Data distribution shifts in time series vital signals (10 time-steps.) Mortality prediction (COVID-19 (Yan et al. 2020) datasets), (b) Sepsis situation identification (SEPSIS (Seymour et al. 2017) dataset)
Li and Jha (2023) Distributional-drift (2 Ep.) Disease classification on physiological signals: for each of the 3 datasets (CovidDeep(Hassantabar et al. 2021), DiabDeep(Yin et al. 2019), and MHDeep(Hassantabar et al. 2022)), 2 episodes were curated by splitting patient into 2 groups.
Chen et al. (2023) Multi-site, multi-vendor (3 Ep., 3 Ep., 6 Ep.) 3 separate applications including (a) Optic disc, (b) cardiac, (c) prostate segmentation were considered in domain shift condition
Li et al. (2024) Multi-vendor (5 Ep.) MR Cardiac segmentation (LV, RV, MYO) with 2 datasets (ACDC and M&M). Episode-1 is from ACDC dataset which is collected from Siemens scanners and the next 4 episodes are from M&M dataset which is collected from 4 vendors (Siemens, Philips, General Electric, and Cannon).
Kim et al. (2024) Multi-site (4 Ep.) Arrhythmia detection on ECG datasets: 4 datasets ((Zheng et al. 2020), (Wagner et al. 2020), (Alday et al. 2020), (Liu et al. 2018)) used as 4 episodes
Aslam et al. (2024) Data distribution shifts in time series vital signals (10 contexts with significant change in mean and standard deviation) Disease outbreak detection via time series signal: 3 datasets (Mpox (Mathieu et al. 2022), Influenza (cdc ), Measles (eur ))) were used separately to curate 10 episodes.
Bayasi et al. (2024) Multi-source (4 Ep.) Skin lesion classification: 4 publicly available skin disease classification datasets (HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018)) were used as 4 episodes, all with same 4 classes of skin disease
Zhu et al. (2024b) Multi-acquisition, multi-equipment (a) Prostate segmentation (6 Ep.), (b) Cardiac segmentation (3 Ep.)
Thandiackal et al. (2024) Multi-source (3 Ep.) Histopathology tissue classification: 3 different datasets (K-19 (Kather et al. 2016), K-16 (Kather et al. 2019), CRC-TP (Javed et al. 2020)) are considered as 3 episodes which contain 7 medically relevant patch classes from H&E stained WSIs of colorectal biopsies
@@ 3.4 Domain incremental scenarios
Domain Incremental Scenario (DIS) is the most popular and frequently observed category of CL scenarios for medical applications. Similar to IIS, here also, the task remains the same over time. However, in contrast to IIS, where data arrives from a single domain, here, the episodic data originates from a different domain or context. This scenario aligns with the idea of learning in a changing environment where the datasets from different domains (e.g., research sites, hospitals, imaging modality, image acquisition protocol, etc.) are incrementally encountered over time, which thus involve covariance shift-induced discrepancies in data.
@@
Yang et al. (2023c) MedMNIST (Yang et al. 2023a) based disease classification (4 Ep.) CIS+DIS: 3 datasets (PathMNIST, DermaMNIST, and OrganAMNIST) as source domain and then 3 more datasets (RetinaMNIST, BreastMNIST, BloodMNIST) as domain shift but classes are added in the class incremental fashion with 1 class at a time
@@
Ceccon et al. (2024b) Chest X-ray based disease classification (7 Ep.) Authors curate a novel scenario termed NIC by interleaving new classes and new instances (from a new dataset as a new domain) in a sequence. The first episode contains a fixed set of classes from ChestX-ray14 (Wang et al. 2017) dataset and then the same classes from CheXpert (Irvin et al. 2019) dataset were considered as second episode. Then some new disjoint classes from ChestX-ray14 were considered in the third and fourth episodes. Thus, authors interleave between new classes and new domains to curate a hybrid scenario containing a total of 7 episodes.
Bayasi et al. (2024) Skin lesion classification (5 Ep.) CIS+DIS: 5 skin lesion classification datasets ( HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), UDA (Codella et al. 2018), BCN (Combalia et al. 2019), PH2 (Mendonça et al. 2013)) with different number of classes (overlapping) were sequentially arranged as 5 episodes which exhibit incremental class and domain simultaneously.
@@ 3.5 Simulated or Hybrid Settings
It may not always be the case that the model observes a particular kind of incremental scenario, such as a domain incremental condition. The CL model developed for handling domain incremental conditions may not perform well for a class incremental condition. Thus, there are attempts to evaluate the same model on different CL scenarios; however, a separate evaluation is followed for each kind of incremental CL scenario. In contrast, having a set of episodes with a mix of CL scenarios is close to real-life conditions. Thus it becomes important to also test the performance in hybrid settings. Table 9 provides a comprehensive overview of works that designed hybrid incremental scenarios.
An attempt in direction is made by Sadafi et al. (2023) where other than pure domain incremental and class incremental, a hybrid incremental scenario is also considered. The authors create a sequence of episodes with both class and domain-incremental cases on multi-site WBC classification datasets. They create a long sequence of episodes with a total of 12 episodes. First, they set up a class incremental scenario within the Matek-19 Matek et al. (2019) dataset by adding 3-4 classes in each subsequent episode up to a total of 4 episodes. After this a new domain, i.e., the INT-20 dataset was considered in the same fashion, contributing 4 more episodes. Finally, a third domain i.e., a new dataset Acevedo-20 Acevedo et al. (2020) was introduced with 4 episodes in class incremental fashion.
Compared to the natural image domain, novel class appearance along with domain shifts is more frequent in medical applications owing to inherent heterogeneity in staining agents, protocols, imaging techniques, vendors, etc. Therefore, there has been recent interest and development of a novel CL scenario that consider new classes as well as new domains/instances, also termed as ‘New Instances and New Classes’ (NIC) scenario Ceccon et al. (2024b). The NIC scenario is particularly relevant for medical image analysis as it addresses the simultaneous occurrence of new types of medical conditions (new classes) and new patient data (new instances), which is common in clinical practice. This scenario emphasizes the need for CL methods that effectively handle class and domain incremental learning. In contrast to traditional class-incremental settings, which often do not account for data shift, Liu et al. (2023a) propose a method for brain tumor segmentation that incrementally adds structures under varying domain shift conditions, such as distinct sites, scanners, or MRI modalities. This approach addresses class-incremental learning by incorporating domain-specific variations. Bayasi et al. (2024) tackle a hybrid continual learning (CL) scenario using five skin lesion datasets from different sources and domains, each containing overlapping and varying numbers of classes. By treating these five datasets as five sequential episodes, their approach exemplifies the integration of both class-incremental and domain-incremental challenges. Similarly, Ceccon et al. (2024b) addresses a hybrid CL scenario in the context of chest X-ray-based disease classification. They interleave disease classes and domains by utilizing datasets from two hospitals, each containing 18 classes, to curate a total of seven episodes. Each episode introduces either new disease classes or a new domain, thus combining class-incremental and domain-incremental learning within their methodology. Medical applications frequently encounter novel classes and domain shifts due to the diversity in staining agents, imaging protocols, and vendor-specific techniques. This has led to the development of new CL scenarios or hybrid learning scenarios.
@@
Further, there are some attempts to curate shifts in datasets and various incremental scenarios instead of collecting real datasets reflecting the situation. For example, Kaustaban et al. (2022) simulated domain shift in H&E stain exhibiting distinct appearance caused by different staining protocols. This dataset is designed to simulate real-world scenarios, where data shifts occur due to differences in scanners, stainers, reagents, and other factors. They considered 9 classes from CRC Kather et al. (2019) which is a H&E stain-based colorectal cancer classification dataset. Each class is divided into 5 disjoint sets for 5 episodes. The first set, regarded as Domain1 is without any alteration and the next 4 sets for each class undergo various changes to create domain shift. A short description of 4 domain shifts is as follows: Domain 2 (increased stain intensity, simulating, for example, concentration increase of the eosin and/or hematoxylin solutions, each with a different extent of change). Domain 3 (decreased eosin stain intensity, simulating, for example, slides prepared from many years ago with fading stain). Domain 4 (change of hue, simulating, for example, change of reagent manufacturer, scanner or stainer) Domain 5 (change of hue and saturation, simulating, for example, change of reagent manufacturer, scanner or stainer). Finally, authors used this augmented dataset for setting class-incremental, domain-incremental, data-incremental, as well as task-incremental scenarios.
@@
Figure 6: The proposed CL taxonomy categorizes various CL models according to their underlying design principles. The abbreviations used are as follows: ER for Experience-replay, GR for Generative-replay, LR for Latent-replay, PF for Prior-focused, DF for Data-focused, and MBO for Memory Buffer Optimization. 1. Hofmanninger et al. (2020a), 2. Perkonigg et al. (2021), 3. Perkonigg et al. (2022), 4. Karthik et al. (2022), 5. Bera et al. (2023), 6. Huang et al. (2023), 7. Wei et al. (2023), 8. Sun et al. (2023b), 9. Zhu et al. (2024a), 10. Li et al. (2022), 11. Byun et al. (2023) 12. Wang et al. (2023b) 13. Thandiackal et al. (2024) 14. Srivastava et al. (2021) 15. Xiao et al. (2023) 16. Bringas et al. (2024) 17. Li and Jha (2023) 18. Baweja et al. (2018), 19. McClure et al. (2018), 20. van Garderen et al. (2019), 21. Zhang and Wang (2019b) 22. Özgün et al. (2020), 23. Zhang et al. (2021), 24. Zhang et al. (2023d), 25. Chen and Tang (2022), 26. Aslam et al. (2024) 27. Zhu et al. (2023), 28. Liu et al. (2023a), 29. Bai et al. (2023) 30. Zhu et al. (2024b) 31. Wu et al. (2024) 32. Karani et al. (2018), 33. González et al. (2022), 34. Yang et al. (2023b) 35. Verma et al. (2024) 36. Qazi et al. (2024) 37. Memmel et al. (2021), 38. Bayasi et al. (2021), 39. Ozdemir et al. (2018), 40. Ozdemir and Goksel (2019), 41. Chee et al. (2023), 42. Li et al. (2020a), 43. Ravishankar et al. (2019), 44. Liu et al. (2022), 45. Akundi and Sivaswamy (2022), 46. Zhang et al. (2023b) 47. Ji et al. (2023) 48. Shu et al. (2022) 49. Zhang et al. (2023c), 50. Sadafi et al. (2023), 51. Hua et al. (2023) 52. Sun et al. (2023a) 53. Ceccon et al. (2024b) 54. Ye et al. (2024) 55. Yang et al. (2023c), 56. Chee et al. (2023), 57. Li et al. (2024), 58. Bayasi et al. (2024)